# Approaching the thermodynamic limit of a bounded one-component plasma

### v1

@@ III.2 Simulation results
Determination of the energies $u_{p\mathrm{e}x}(\Gamma,N)$ and $u_{b\mathrm{e}x}(\Gamma,N)$ makes it possible to obtain the ionic compressibility factor (16) In Fig. 8, it is compared with the same quantity estimated from the virial (24). It is noteworthy that (i) estimates from the energies are in a good agreement with those from virial, however, the latter have much larger scatter and (ii) within the accuracy attained in our simulation, $Z_{i}$ seems to be size-independent. Such peculiarities hold true irrespective of $\Gamma$. Thus, we take the average of all $Z_{i}$ available for different $N$ at the same $\Gamma$ as an estimate of its thermodynamic limit, the error being calculated by application of a standard statistical method to the sample of the size up to 9. The resulting dependence $Z_{i}(\Gamma)$ is shown in Fig. 9. As is seen, the results from energy and virial are hardly distinguishable, except for the highest $\Gamma$, which validates correctness of $Z_{i}$ determination. At $\Gamma<0.3$, the obtained $Z_{i}$ almost coincides with the ionic compressibility factor $1-{\Gamma^{3/2}}/2\sqrt{3}$ that results from the Debye–Hückel approximation, as it must. In the interval $0.1\leq\Gamma\leq 30$, $Z_{i}$ can be approximated by the function
${Z_{i}}(\Gamma)={c_{1}}+{c_{2}}\tanh\left({\displaystyle{{{c_{3}}-\ln\Gamma}\over{{c_{4}}}}}\right),$ (43)
with $c_{1}=0.4613762483$, $c_{2}=0.5260100823$, $c_{3}=1.4830111629$, and $c_{4}=1.3508409652$. Note that at $\Gamma>30$, $Z_{i}\simeq 0$ as it follows from Eq. (13), and in the entire range of $\Gamma$, $Z_{i}>0$. Note that at $\Gamma>300$, images of the cold ions from a spherical boundary are so rare that they can be neglected. Therefore, the difference between BOCP and Coulomb cluster diminishes, and for these systems, $Z_{i}$ must coincide. In [62], a negative value $Z_{i}=-0.04\pm 0.06$ was obtained at $\Gamma=500$ as a result of the substantial simulation inaccuracy.
Figure 9: Ionic compressibility factor $Z_{i}$ determined from MD simulation using Eqs. (16) (circles) and (24) (squares). Dashed curve indicates the approximation (43) of $Z_{i}$ from energies; solid line, $Z_{i}$ for the entire system (21); and the dashed-dotted line, $Z_{i}$ in the Debye–Hückel apprximation.
@@ IV.1 Simulation procedure
A common way to calculate properties of an OCP system is to use the Ewald potential, which is typically split into two parts: $\text{erfc}(\alpha r)/r$, dominant at short and medium distances within the simulation cell, and $\text{erf}(\alpha r)/r$, a smooth, long-range function treated in reciprocal space. Here, $\alpha$ is the Ewald damping (splitting) parameter, chosen such that the real-space contribution at the cutoff length $r_{c}$ satisfies the condition $\text{erfc}(\alpha r_{c})/r_{c}\lesssim\varepsilon_{0}$, where $\varepsilon_{0}$ is the target relative error for forces. Interactions within $r_{c}$ are calculated directly in real space, while interactions beyond $r_{c}$ are included via the reciprocal-space sum over periodic images. It is assumed that thermodynamic properties of the system are effectively independent of the precise cutoff, as long as $\alpha$ and $r_{c}$ are chosen to meet the desired accuracy $\varepsilon_{0}$.
To examine the effect of the cutoff length on derived thermodynamic quantities arising from specific computational algorithms, we performed a series of MD simulations using the stable release of LAMMPS (29 August 2024) [69], compiled with the ScaFaCoS library v1.0.4 [70].
The total number of particles $N$ was fixed within each simulation, but varied between different simulations and the exact values will be specified later. Periodic boundary conditions were applied in all three directions of a cubic cell with the edge length $L=(4\pi N/3)^{1/3}$. The integration time step was chosen to be dependent on the Coulomb coupling parameter $\Gamma$ as $\tau=10^{-3}(0.2\Gamma+2)$.
Ions of mass $m=1$ bearing the charge $z=1$ and atom style charge were initialized randomly for fluid simulations and as a bcc lattice for solid-state simulations. Initial velocities were assigned the Maxwellian distribution corresponding to temperature $T=1/\Gamma$. In both cases, the pair style coul/long was used with the cutoff length $r_{c}$. Long-range interactions were handled with the $k$-space solver ewald for simulations employing the Ewald potential [58], or with pppm for simulations using the PPPM method [67], in both cases with the target relative accuracy $\varepsilon_{0}=10^{-6}$.
@@
IV.2 Effect of the ionic pressure
@@
In order to demonstrate the dependence of ionic compressibility factor of a system on chosen cutoff length with fixed relative error, a series of MD simulations were performed under simulation conditions specified in Sec. IV.1, while the total number of ions was chosen to be $N=10^{3}$ and $k$-space style pppm was used.
@@
Figure 10: Ionic compressibility factor $Z_{i}$ for OCP as a function of the LAMMPS cutoff length $r_{c}$ for the cases of random initialization (fluid) at $\Gamma=0.1$ (circles), $\Gamma=10$ (squares), $\Gamma=160$ (diamonds), and $\Gamma=200$ (triangles up) and initialization as bcc (solid) at $\Gamma=160$ (triangles down) and $\Gamma=200$ (stars). Dashed and dotted lines illustrate $Z_{i}$ obtained for BOCP at $\Gamma=0.1$ and $200$, respectively.
@@
The ion compressibility factor $Z_{i}$ shows a strong dependence on the cutoff length $r_{c}$, as illustrated in Fig. 10, and this dependence becomes stronger with the increase in $\Gamma$ (cf. [63]). In all cases, $Z_{i}$ grows as $r_{c}$ increases, indicating that the virial is also sensitive to $r_{c}$. For weak coupling ($\Gamma=0.1$), $Z_{i}$ varies only slightly with $r_{c}$ and stays close to the BOCP reference value, demonstrating that the effect is present but relatively modest. At intermediate coupling ($\Gamma=10$), the growth of $Z_{i}$ with $r_{c}$ becomes clearly visible though still moderate. In the strongly coupled regimes ($\Gamma=160$ and $\Gamma=200$), the dependence turns dramatic: $Z_{i}$ rises by orders of magnitude as $r_{c}$ increases. Thus, ionic compressibility factor increases with increasing cutoff distance regardless of coupling parameter, however while it is weak but detectable at small $\Gamma$, the dependency strengthens steadily at intermediate coupling, and dominates the behavior at strong coupling. The choice of initial configuration (fluid vs. solid) appears to have no decisive influence on the resulting $Z_{i}$ values.
From a physical perspective, a strong dependence of $Z_{i}$ on $r_{c}$ at high $\Gamma$ reflects the emergence of interfacial effects in condensed states. In this regime, truncation of the Coulomb potential effectively modifies the balance between bulk and surface terms of the virial. The rapid increase of $Z_{i}$ with $r_{c}$ can therefore be interpreted as an artifact of introducing a spurious surface tension in a nominally homogeneous system with periodic boundaries. While relatively small in weakly coupled plasmas, this artificial surface term dominates the virial at strong coupling, leading to nonphysical values of $Z_{i}$ depending on $r_{c}$.
@@
Figure 11: LAMMPS cutoff length $r_{c}$ ensuring equality of the compressibility factor calculated from the virial for OCP and the ionic compressibility factor of BOCP (43) for the Ewald and PPPM calculation methods (circles and diamonds, respectively), and equality of the compressibility factor calculated from the OCP virial and the OCP potential energy (21) for the Ewald and PPPM calculations (squares and stars, respectively). Solid line is the dependence $r_{c}(\Gamma)$ (44) and (45) fitting dots.
@@
In this context, it becomes necessary to determine the proper cutoff length $r_{c}$ that yields the correct ionic pressure. This was done by numerically solving the equation $Z_{\mathrm{MD}}(r_{c})=Z_{i}(\Gamma)$ at the same $\Gamma$ for $r_{c}$ , i.e., by varying $r_{c}$ in the range from 0.1 to 16 and comparing the resulting $Z_{\mathrm{MD}}$ with the BOCP reference value $Z_{i}(\Gamma)$ (43). In all simulations, the total number of ions was fixed at $N=29^{3}$, and the ions were initialized randomly, except for the case $\Gamma=300$, where the bcc lattice initialization was used. We also attempted to determine the cutoff $r_{c}$ that ensures the correct compressibility factor of the entire system, by solving $Z_{\mathrm{MD}}(r_{c})=Z$ with $Z$ defined by (21). The resulting dependence of the optimal cutoff length $r_{c}$ on $\Gamma$ is shown in Fig. 11. For the ionic compressibility factor criterion, the cutoff obtained with the Ewald method exhibits a smooth monotonic increase with $\Gamma$ approaching $r_{c}\approx 3$ at strong coupling. The PPPM method follows the same trend and shows no significant deviation from the Ewald results, while reducing the computational cost. The fitted curve is given by the expressions
@@
where $\eta=\ln(\Gamma)$, $a_{0}=1.616474$, $a_{1}=0.205984$, $a_{2}=0.0373460$, $a_{3}=0.0269303$, and $a_{4}=0.00775118$ at $0.1\leq\Gamma\leq 5$, and
@@
The compressibility factor of the entire system, which includes the neutralizing background, is lower than the ionic one, and therefore requires a significantly smaller cutoff length. However, for $r_{c}<1$, the number of reciprocal lattice vectors required for the long-range (reciprocal space) part, $k_{c}\sim\ln(\varepsilon_{0}^{-2})/r_{c}$, increases rapidly, leading to excessive memory usage. This memory limitation is the main reason for the unclear results obtained when attempting to determine the proper cutoff length for the full-system compressibility factor, as well as for the large uncertainty in the region $\Gamma<1$. Thus, the uncertainty shown in Fig. 11 is calculated as a difference between the approximation of $r_{c}$ in the last successful run and the previous approximation. Such failure in $r_{c}$ determination confirms the main conclusion made in Sec. II.2 that it is the ionic compressibility factor that is found by the virial from LAMMPS calculations rather than a full compressibility factor of the entire system.
@@
In order to explore other possible effects of the cutoff length, the fluid–solid phase transition in the OCP was studied at $\Gamma=10\mbox{--}250$ with a step of 10. At each $\Gamma$, four simulations were performed: (i) $r_{c}=14$ with random (fluid) initialization, (ii) $r_{c}=14$ with the bcc lattice initialization, (iii) $r_{c}=r_{c}(\Gamma)$ with random initialization, where $r_{c}(\Gamma)$ is calculated using Eqs. (44) and (45), and (iv) $r_{c}=r_{c}(\Gamma)$ with the bcc lattice initialization. In all four cases, the $k$-space style ewald was employed. For simulations starting from the fluid phase, the number of ions was set to $29^{3}$, while for those initialized as a bcc lattice, $N=10^{3}$ was used.
@@
The thermal fraction of the potential energy calculated from the system potential energy as $\Gamma(u-u_{0})$ is shown in Fig. 12, where uncertainties are calculated as a difference in the averaged results between simulations performed under the same conditions but using different initial random number generator seeds. All simulation protocols reproduce the known dependence of this quantity, closely following the MC results by Slattery et al. [43]. The choice of different cutoff lengths ($r_{c}=14$ or $r_{c}(\Gamma)$) has a little effect, and both the fluid and bcc initializations converge to almost identical values in a single-phase state. This demonstrates that in the single-phase regime, the system equilibrates effectively, regardless of the initialization type and cutoff treatment. In contrast, differences become apparent near the fluid–solid transition. With the optimized cutoff $r_{c}(\Gamma)$, simulations initialized in the fluid state exhibit a sharp transition at $\Gamma=\mbox{180--200}$, while runs with fixed $r_{c}=14$ display the transition shifted to lower $\Gamma$, though following the same trend. For the bcc initialization, the system with $r_{c}=14$ undergoes the transition near $\Gamma\approx 150$, whereas those using the optimized $r_{c}(\Gamma)$ show the transition delayed to higher $\Gamma$. Overall, the metastable region of the OCP proves to be narrower when a fixed cutoff is used, while the optimized $r_{c}(\Gamma)$ yields significantly broader metastability window. Such behavior can be rationalized in terms of the dependence of the compressibility factor on the cutoff length. Smaller cutoff values reduce the virial contribution, thereby decreasing the effective surface tension. A lower surface tension reduces the Gibbs free energy barrier for nucleation $\Delta G\sim\sigma^{3}$ and enhances the nucleation rate $J\sim\exp(-\Delta G/T)$ thus making the emergence of phase transition earlier.
V Conclusion

### v2

@@ III.2 Simulation results
Determination of the energies $u_{p\mathrm{e}x}(\Gamma,N)$ and $u_{b\mathrm{e}x}(\Gamma,N)$ allows one to obtain the ionic compressibility factor (16). In Fig. 8, it is compared with the same quantity estimated from the virial (24). It is noteworthy that (i) estimates from the energies are in a good agreement with those from the virial, however, the latter have much larger scatter and (ii) within the accuracy attained in our simulation, $Z_{i}$ seems to be size-independent. Such peculiarities hold true irrespective of $\Gamma$. Thus, we take the average of all $Z_{i}$ available for different NN at the same $\Gamma$ as an estimate of its thermodynamic limit, the error being calculated by application of a standard statistical method to the sample of size up to nine, by the number of BOCP sizes used. The resulting dependence $Z_{i}(\Gamma)$ is shown in Fig. 9 and listed in Table 2. As is seen, the results from energy and virial are hardly distinguishable, except for the highest $\Gamma$, which validates the overall correctness of $Z_{i}$ determination. At $\Gamma<0.3$, the obtained $Z_{i}$ almost coincides with the ionic compressibility factor $Z=1-{\Gamma^{3/2}}/2\sqrt{3}$ that results from the Debye–Hückel approximation, as it must. In the interval $0.1\leq\Gamma\leq 27.72$, $Z_{i}$ can be approximated by the function
${Z_{i}}(\Gamma)={c_{1}}+{c_{2}}\tanh\left({\displaystyle{{{c_{3}}-\ln\Gamma}\over{{c_{4}}}}}\right)$ (43)
with $c_{1}=0.461376$, $c_{2}=0.526010$, $c_{3}=1.483011$, and $c_{4}=1.350840$. Note that at $\Gamma>27.72$, $Z_{i}\simeq 0$ as it follows from Eq. (13), and in the entire range of $\Gamma$, $Z_{i}>0$. It is worth mentioning that at $\Gamma>300$, reflections of the cold ions from a spherical boundary are so rare that they can be neglected. Therefore, the difference between BOCP and Coulomb cluster diminishes, and for these systems, $Z_{i}$ must coincide. In [73], a negative value $Z_{i}=-0.04\pm 0.06$ was obtained at $\Gamma=500$ as a result of the substantial simulation inaccuracy.
Figure 9: Ionic compressibility factor $Z_{i}$ determined from MD simulation using Eqs. (16) (circles) and (24) (squares). Dashed curve indicates the approximation (43) of $Z_{i}$ from the energies; solid line, the compressibility factor $Z$ for the entire system (21); and the dashed-dotted line, $Z$ in the Debye–Hückel approximation.
@@ IV.1 Simulation procedure
A common way to calculate properties of an OCP system is to use the Ewald potential, which is typically split into two parts: $\text{erfc}(\alpha r)/r$, dominant at short and medium distances within the simulation cell, and $\text{erf}(\alpha r)/r$, a smooth, long-range function treated in reciprocal space. Here, $\alpha$ is the Ewald damping (splitting) parameter, chosen such that the real-space contribution at the cutoff length $r_{c}$ satisfies the condition $\text{erfc}(\alpha r_{c})/r_{c}\lesssim\varepsilon_{0}$, where $\varepsilon_{0}$ is the target relative error for forces. Interactions within $r_{c}$ are calculated directly in real space, while interactions beyond $r_{c}$ are included via the reciprocal-space sum over periodic images. It is assumed that thermodynamic properties of the system are essentially independent of the precise cutoff, as long as $\alpha$ and $r_{c}$ are chosen to meet the desired accuracy $\varepsilon_{0}$.
To examine the effect of the cutoff length on derived thermodynamic quantities arising from specific computational algorithms, we performed a series of MD simulations using the stable release of LAMMPS (Large-scale Atomic/Molecular Massively Parallel Simulator, 29 August 2024) [80] compiled with the ScaFaCoS library v1.0.4 [81].
The total number of particles NN was fixed within each simulation, but varied between different simulations and the exact values will be specified later. Periodic boundary conditions were applied in all three directions of a cubic cell with the edge length $L=(4\pi N/3)^{1/3}$. The integration time step was chosen to be dependent on the Coulomb coupling parameter $\Gamma$ as $\tau=10^{-3}(0.2\Gamma+2)$.
Ions of mass $m=1$ bearing the charge $ze=1$ were initialized randomly for fluid simulations and as a bcc lattice for solid-state simulations. Initial velocities were assigned the Maxwellian distribution corresponding to temperature $T=1/\Gamma$. In both cases, the energies and forces were calculated by a direct summation in the real space inside a sphere of the radius $r_{c}$ around each ion and by summation in $k$-space outside the sphere. Thus, long-range interactions were handled with the $k$-space Ewald solver for simulations employing the Ewald method [70], or with PPPM solver for simulations using the PPPM method [78]; in both cases, with the target relative accuracy $\varepsilon_{0}=10^{-6}$. The Ewald method implies a direct calculation in the $k$-space, while the PPPM method assigns charges to a finely-spaced mesh in the simulation box. Then the Poisson equation for the electrostatic potential is solved in the $k$-space, and the force at each mesh point is calculated by numerically differentiating this potential. The force acting on an ion is evaluated as an interpolation of the mesh forces.
@@
IV.2 Correction to the calculation of ionic pressure
@@
In order to demonstrate the dependence of the ionic compressibility factor of a system on the chosen cutoff length at fixed relative error, a series of MD simulations were performed under simulation conditions specified in Sec. IV.1, while the total number of ions was chosen to be $N=10^{3}$ and the $k$-space PPPM method was used.
@@
Figure 10: Ionic compressibility factor $Z_{i}$ for OCP as a function of the LAMMPS cutoff length $r_{c}$ for the cases of random initialization (fluid) at $\Gamma=0.1$ (circles), $\Gamma=10$ (squares), $\Gamma=160$ (diamonds), and $\Gamma=200$ (triangles up) and initialization as bcc (solid) at $\Gamma=160$ (triangles down) and $\Gamma=200$ (stars). Dashed and dotted lines show $Z_{i}$ obtained for BOCP at $\Gamma=0.1$ and $200$, respectively.
@@
The ion compressibility factor $Z_{i}$ (24) shows a strong dependence on the cutoff length $r_{c}$, as illustrated in Fig. 10 (a vertical shift of $Z_{i}$ by more than $14$ units is necessary to display it on a logarithmic scale), and this dependence becomes stronger with the increase in $\Gamma$ (cf. [74]). In all cases, $Z_{i}$ grows as $r_{c}$ increases, indicating that the virial is also sensitive to $r_{c}$. For weak coupling ($\Gamma=0.1$), $Z_{i}$ varies only slightly with $r_{c}$ and stays close to the BOCP reference value, demonstrating that the effect is present but relatively modest. At intermediate coupling ($\Gamma=10$), the growth of $Z_{i}$ with $r_{c}$ becomes clearly visible though still moderate. In the strongly coupled regimes ($\Gamma=160$ and $\Gamma=200$), the dependence turns dramatic: $Z_{i}$ rises by orders of magnitude as $r_{c}$ increases. The choice of initial configuration (fluid vs. solid) appears to have no decisive influence on the resulting $Z_{i}$ values. A strong dependence of $Z_{i}$ on $r_{c}$ at high $\Gamma$ may be indicative of some artifact resulting, e.g., from a discontinuity of the effective interionic potential (and/or its spatial derivatives) at $r=r_{c}$ and the consequent singularity of its derivatives in coordinates at this point.
@@
Figure 11: LAMMPS cutoff length $r_{c}$ ensuring equality of the compressibility factor calculated from the virial for OCP and the ionic compressibility factor of BOCP (43) for the Ewald and PPPM calculation methods (circles and diamonds, respectively), and equality of the compressibility factor calculated from the OCP virial and the OCP electrostatic energy (21) for the Ewald and PPPM methods (squares and stars, respectively). Solid line is the dependence $r_{c}(\Gamma)$ (44) and (45) fitting circles.
@@
In this context, it becomes necessary to determine the proper cutoff length $r_{c}$ that yields the correct ionic pressure. This was done by numerically solving the equation $Z_{\mathrm{MD}}(r_{c})=Z_{i}(\Gamma)$ for $r_{c}$ at the same $\Gamma$, i.e., by varying $r_{c}$ in the range from 0.1 to 16 and comparing the resulting $Z_{\mathrm{MD}}$ with the BOCP reference value $Z_{i}(\Gamma)$ (43). In all simulations, the total number of ions was fixed at $N=29^{3}$, and the ions were initialized randomly, except for the case $\Gamma=300$, where the bcc lattice initialization was used. We also attempted to determine the cutoff $r_{c}$ that ensures the correct compressibility factor of the entire system, by solving $Z_{\mathrm{MD}}(r_{c})=Z(\Gamma)$ with $Z(\Gamma)$ defined by (21).
The resulting dependence of the optimal cutoff length $r_{c}$ on $\Gamma$ is shown in Fig. 11. For the ionic compressibility factor criterion, the cutoff obtained with the Ewald method exhibits a smooth monotonic increase with $\Gamma$, with the cutoff approaching $r_{c}\simeq 3$ at strong coupling. The PPPM method follows the same trend and shows no significant deviation from the Ewald results and also reduces the computational cost. The fitted curve is given by the expressions
@@
where $\eta=\ln(\Gamma)$, $a_{0}=1.61637$, $a_{1}=0.21068$, $a_{2}=0.035588$, $a_{3}=0.024795$, and $a_{4}=0.0078812$ at $0.1\leq\Gamma\leq 4$, and
@@
The compressibility factor of the entire system, which includes the neutralizing background, is lower than the ionic one, and therefore requires a significantly smaller cutoff length. However, for $r_{c}<1$, the number of reciprocal lattice vectors required for the long-range part in reciprocal space, which is proportional to $\left[\ln(1/\varepsilon_{0})/r_{c}\right]^{3}N$, increases rapidly, leading to excessive memory usage. This memory limitation is the main source of large uncertainties obtained when attempting to determine the proper cutoff length for the full-system compressibility factor $Z$. The uncertainties shown in Fig. 11 are calculated as the difference between the approximation of $r_{c}$ in the last successful run and the previous approximation. A failure in the determination of $r_{c}$ in the latter case confirms the main conclusion made in Sec. II.2 that it is the ionic compressibility factor that is obtained using the virial from LAMMPS calculations rather than the compressibility factor for the entire system.
V Ambiguity of force-based quantities for the Coulomb OCP
An explicit non-physical dependence of the ionic pressure on $r_{c}$ naturally leads to the hypothesis that existing convergence-acceleration schemes for the Coulomb sums reproduce the excess energy with high accuracy, but introduce non-negligible ambiguous distortions in the ionic forces. Once this force-level ambiguity is admitted, its consequences for kinetic and interfacial properties follow directly. For the fluid and solid OCP, the phonon spectra are determined by the dynamical matrix
$D_{\alpha\beta}(i,j)=\displaystyle{\partial^{2}u\over\partial x_{i,\alpha}\partial x_{j,\beta}},$ (46)
where $x_{i,\alpha}$ is the $\alpha$ component of the coordinate of $i$th ion, i.e., by the second derivatives of potential, equivalently, by derivatives of the effective forces acting between the $i$th and $j$th ions. Even a small, systematic distortion of these forces may lead to a considerable error in the phonon frequencies, the group velocities, and, generally, in any quantity governed by the lattice dynamics. Similarly, the Green–Kubo expressions for transport coefficients involve microscopic fluxes that depend explicitly on forces. For example, the potential term of the virial stress tensor entering the Green–Kubo formulas for shear viscosity and thermal conductivity has the form
$\sigma^{\mathrm{pot}}_{\alpha\beta}=\displaystyle{1\over 2V}\sum_{i\neq j}r_{ij,\alpha}f_{ij,\beta}.$ (47)
Even if the effective pair interionic potential reproduces energies to high accuracy, any systematic bias in forces leads to biased Green–Kubo integrals for the viscosity, electrical conductivity, and thermal conductivity.
The situation is even more delicate for interfacial properties, such as the surface tension $\gamma$, which is expressed in terms of the stress tensor $\sigma_{\alpha\beta}$ as
$\gamma=\int_{-\infty}^{\infty}\left[\sigma_{nn}(z)-\sigma_{nn}^{\mathrm{bulk}}\right]dz,$ (48)
where $\sigma_{nn}^{\mathrm{bulk}}$ is the stress tensor component in the bulk phase and $z$ is the coordinate normal to the interface. In the classical OCP, for which no liquid–solid binodal exists, surface tension manifests itself only in metastable states involving crystallites in the fluid or fluid inclusions in the crystal. In such situations, $\gamma$ is determined entirely by subtle differences between the stress components in an inhomogeneous region, and it is therefore extremely sensitive to long-range force errors, even when the bulk energy is estimated almost perfectly. Since the Gibbs free energy barrier for the fluid–solid nucleation $\Delta G\sim\gamma$, the nucleation rate $J\sim\exp(-\Delta G/T)$ is very sensitive to $\gamma$. This, in turn, can result in a pronounced correlation between the width of a fluid–solid metastable region observed in MD simulations and $\gamma$, i.e., $r_{c}$.
@@ V Ambiguity of force-based quantities for the Coulomb OCP
In order to check this, the fluid–solid phase transition in the OCP was studied at $\Gamma=10\mbox{--}250$ with a step of 10. At each $\Gamma$, four simulations were performed: (i) $r_{c}=14$ with random (fluid) initialization, (ii) $r_{c}=14$ with the bcc lattice initialization, (iii) $r_{c}=r_{c}(\Gamma)$ with random initialization, where $r_{c}(\Gamma)$ is calculated using Eqs. (44) and (45), and (iv) $r_{c}=r_{c}(\Gamma)$ with the bcc lattice initialization. For simulations starting from the fluid phase, the number of ions was set to $29^{3}$, while for those initialized as a bcc lattice, $N=10^{3}$ was used.
@@
The thermal fraction of the potential energy calculated from the system potential energy as $\Gamma(u-u_{0})$ is shown in Fig. 12, where uncertainties are calculated as the difference between the averaged results of simulations performed under the same conditions but with different initial seeds of the random number generator. In a single-phase state, all simulation protocols reproduce the known dependence of $\Gamma(u-u_{0})$ closely following the MC results by Slattery et al. [54]. The choice of different cutoff lengths ($r_{c}=14$ or $r_{c}(\Gamma)$) has little effect, and both the fluid and bcc initializations converge to almost identical values. This demonstrates that in the single-phase regime, the system equilibrates effectively, regardless of the initialization type and cutoff treatment. In contrast, differences become apparent near the fluid–solid transition. With the optimized cutoff $r_{c}(\Gamma)$, simulations initialized in the fluid state exhibit a sharp transition at $\Gamma=\mbox{180--200}$, while runs with fixed $r_{c}=14$ display the transition shifted to lower $\Gamma$, though following the same trend. For the bcc initialization, the system with $r_{c}=14$ undergoes the transition near $\Gamma\approx 150$, whereas those using the optimized $r_{c}(\Gamma)$ show the transition delayed to lower $\Gamma$. Overall, the metastable region of the OCP proves to be narrower when a fixed cutoff is used, while the optimized $r_{c}(\Gamma)$ yields a significantly broader metastability window. Such behavior can be rationalized in terms of the increasing dependence of the compressibility factor on the cutoff length in the vicinity of the optimized $r_{c}(\Gamma)$. Since $\sum\nolimits_{i=1}^{N}{\left\langle{{{\mathbf{r}}_{i}}{{\mathbf{f}}_{i}}}\right\rangle}<0$, it is seen from (24) and Fig. 10 that a larger $r_{c}$ reduces the virial modulus thereby decreasing the surface tension $\gamma$ (48). A lower surface tension reduces the Gibbs free energy barrier for nucleation and enhances the nucleation rate thus accelerating the onset of phase transition and leading to the narrower metastable region at $r_{c}=14$, in accordance with our simulations.
Our findings indicate that standard Ewald-type treatments correspond to an OCP-like model, whose properties depend sensitively on how the calculation acceleration of the long-range Coulomb sums is controlled numerically, rather than the classical OCP. At the formal level, the exact Ewald summation corresponds to a unique OCP Hamiltonian and can, in principle, be evaluated to arbitrary precision. In actual MD simulations, however, one has to introduce finite real-space cutoffs, finite $k$-space resolution, and target error tolerances, in addition to the unavoidable finite-size effects. A specific choice of $r_{c}$ may lead to sizeable, systematic distortions in forces and virial-based quantities such as the ionic compressibility factor. Hence, with respect to the kinetic and nonequilibrium properties, we are effectively testing an “Ewald-regularized OCP” with specific realization of the long-range summation algorithm. This motivates the development of methods that define the OCP EOS in terms of energy quantities only. The excess energy can be separated into the particle–particle and particle–background contributions like $u_{p\mathrm{ex}}$ and $u_{b\mathrm{ex}}$ and used to construct the corresponding ionic compressibility factor $Z_{i}^{(E)}(\Gamma)$ as well as related thermodynamic functions without any reference to the force virial. Apparently, this can be made by tuning the real-space cutoff $r_{c}$ or other realization-specific parameters. Then $Z_{i}^{(E)}$ could be used as a stringent consistency condition: only those implementations, for which the ionic compressibility factor obtained from forces matches $Z_{i}^{(E)}$ for the given $\Gamma$, should be trusted for the calculation of phonon spectra, transport coefficients, and interfacial properties.
VI Conclusion